please assist ( anyone please help )
Bush juices (BJ) Lts and Wild Tates (WT) Ltd. Produce and sell natural fruit juices. BJ, at the beginning of October 2006, conducted a study to establish its market share vis – a-vis that of NT, the only other competitor in natural fruit juice market. The study-involved observation of loyalty shifting habits of a total pool of 19200 natural fruit juice consumers, distributed equally between the two companies. Records by the end of the month indicated that 2400 consumers had switched from BJ to WT and half of WT ‘s had shifted to BJ. The total consumer pool is expected to grow to exponentially, at the rate of 10 % per month, up to the end of June 2007 after which it is expected to stabilize with negligible growth thereafter Each consumer of natural fruit juice, generates an average monthly revenue of 400 $ to the selling company.
Assuming that the October 2006 shift pattern persists into the future months, determine the expected revenue generated by each company during
1. November 2006
2. December 2006
3. Any month at equilibrium


Answer by stanbon(75887)

BJ October DATA:
Start: (1/2)19600=9800
Then: 9800-2400+0.5(9800)+0.1(9800)=
End: =13,280 for BJ
Therefore, WT End: 1.1*19200-13280=7840
Total end=1.1*19600=21560
October: Revenue Generated:
BJ Oct. Revenue: 13280*400=$5,312,000
WT Oct. Revenue: 7840*400=$3,136,000
------------
BJ Novenber DATA:
Start: 13,280
Then: 13280-2400+0.5(7840)+0.1*13280=
End: 16128 for BJ
Therefore WT End:21560-16128=5432
November Revenue Gererated:
BJ Nov. Revenue: 16128*400=$6,451,200
WT Nov. Revenue: 5432*400=$2,172,800
---------
I'll leave December to you.
Also, I don't see any equilibrium between
BJ and WT Revenue. BJ's manpower is
increasing and WT's is decreasing as far
as I can see.
--------
Hope this helps.
Cheers,
Stan H.